# On the Genericity of the Spectrum Intervalization for Multi-Frequency Quasiperiodic Schrödinger Operators

### v1

@@ Conclusion.
We prove the lemma using Cartan-type estimates for analytic functions and a parameter-space argument. Let
$V(x;\mathbf{c})=\sum_{m\in\mathbb{Z}^{d}:|m|\leq n}c_{m}e^{2\pi im\cdot x}$
be the trigonometric polynomial potential, where $\mathbf{c}=(c_{m})\in\mathbb{R}^{N}$. For fixed $h\in\mathbb{T}^{d}$, define the analytic functions:
$F_{h}(x;\mathbf{c})=V(x+h;\mathbf{c})-V(x;\mathbf{c}),\quad G_{h}(x;\mathbf{c})=g_{V,h,i,j}(x;\mathbf{c}),$ (3.3)
where $g_{V,h,i,j}$ is the determinant function derived from $V$ as in condition (iii) of class $\mathfrak{G}$. The Cartan estimate requires that for any $K\gg 1$ and $\|h\|\geq\exp(-K)$,
$\text{mes}\left\{x\in\mathbb{T}^{d}:\min\left(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|\right)<\exp(-K)\right\}\leq\exp(-K^{\mathfrak{c}_{1}}),$ (3.4)
where $\mathfrak{c}_{1}>0$ is a constant. Define the bad set for fixed $h$ and $K$ as:
$\mathcal{B}^{(iii)}(h,K)=\left\{\mathbf{c}\in\mathbb{R}^{N}:\text{mes}\left\{x:\min\left(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|\right)<\exp(-K)\right\}>\exp(-K^{\mathfrak{c}_{1}})\right\}.$ (3.5)
We aim to show that the overall bad set $\mathcal{B}^{(iii)}=\bigcup_{h\in\mathbb{T}^{d}}\bigcup_{K\geq K_{0}}\mathcal{B}^{(iii)}(h,K)$ has measure zero.
The proof proceeds in three steps:
(1) Uniform Cartan estimate for bounded parameters: For any fixed $R>0$, consider the ball $B_{R}=\{\mathbf{c}\in\mathbb{R}^{N}:\|\mathbf{c}\|\leq R\}$. Since $F_{h}$ and $G_{h}$ are analytic in $x$ and linear in $\mathbf{c}$, the family $\{F_{h}(\cdot;\mathbf{c}),G_{h}(\cdot;\mathbf{c}):\mathbf{c}\in B_{R}\}$ is uniformly bounded and equicontinuous on $\mathbb{T}^{d}$. By the Cartan-type estimate for analytic functions (see Cartan [7] or Goldstein et al. [11]), there exist constants $C_{R}>0$ and $\delta_{R}>0$ (depending on $R$ and $h$) such that for all $\mathbf{c}\in B_{R}$ with $F_{h}$ and $G_{h}$ not both identically zero,
$\text{mes}\left\{x:\min\left(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|\right)<\epsilon\right\}\leq C_{R}\epsilon^{\delta_{R}}.$ (3.6)
To derive this, recall that for an analytic function $f$ on $\mathbb{T}^{d}$ with analytic norm $\|f\|_{\mathcal{A}}$, the Cartan estimate guarantees that for any $\epsilon>0$,
$\text{mes}\left\{x:|f(x)|<\epsilon\right\}\leq C\exp\left(-\frac{\delta}{\|f\|_{\mathcal{A}}}\log\frac{1}{\epsilon}\right),$ (3.7)
where $C,\delta>0$ depend on the dimension $d$. Applying this to $f=\min(|F_{h}|,|G_{h}|)$ and noting that $\|F_{h}\|_{\mathcal{A}}\leq 2\|V\|_{\mathcal{A}}$ and $\|G_{h}\|_{\mathcal{A}}\leq C^{\prime}\|V\|_{\mathcal{A}}$ for some $C^{\prime}>0$, we obtain (3.6) with $\delta_{R}$ proportional to $1/\|V\|_{\mathcal{A}}$. Since $\|V\|_{\mathcal{A}}$ is uniformly bounded for $\mathbf{c}\in B_{R}$, the constants $C_{R}$ and $\delta_{R}$ are uniform over $B_{R}$.
(2) Measure zero for bounded parameters:. Now, fix $h$ and $R$. The set of $\mathbf{c}\in B_{R}$ for which $F_{h}$ or $G_{h}$ is identically zero is contained in a proper linear subspace of $\mathbb{R}^{N}$. Specifically, if $F_{h}\equiv 0$, then for all $x$,
$\sum_{m}c_{m}\left(e^{2\pi im\cdot(x+h)}-e^{2\pi im\cdot x}\right)=0,$ (3.8)
which implies that $c_{m}(e^{2\pi im\cdot h}-1)=0$ for all $m$. For generic $h$, $e^{2\pi im\cdot h}\neq 1$ for $m\neq 0$, so $c_{m}=0$ for $m\neq 0$. Thus, the set of such $\mathbf{c}$ has measure zero. Similarly for $G_{h}$. Hence, for $\mathbf{c}\in B_{R}$ outside this null set, the Cartan estimate (3.6) applies. Choose $K_{0}(R,h)$ large enough such that for all $K\geq K_{0}(R,h)$, we have
$C_{R}\exp(-\delta_{R}K)\leq\exp(-K^{\mathfrak{c}_{1}}).$ (3.9)
Then for $K\geq K_{0}(R,h)$, the bad set $\mathcal{B}^{(iii)}(h,K)\cap B_{R}$ is empty. For $K<K_{0}(R,h)$, the bad set is contained in the null set where $F_{h}$ or $G_{h}$ is identically zero. Thus, for fixed $h$ and $R$, the set $\bigcup_{K\geq K_{0}}\mathcal{B}^{(iii)}(h,K)\cap B_{R}$ has measure zero.
(3) Union over $h$ and $K$, and unbounded parameters:. To handle the union over $h\in\mathbb{T}^{d}$, we use a covering argument. For each $K\geq K_{0}$, cover $\mathbb{T}^{d}$ with a grid of points $h_{j}$ such that the number of points is $N(K)=O(\exp(Kd))$ and the mesh size is $\exp(-K)$. By the uniform continuity of $F_{h}$ and $G_{h}$ in $h$, for any $h$ in the $\exp(-K)$-neighborhood of $h_{j}$, we have
$|F_{h}(x;\mathbf{c})-F_{h_{j}}(x;\mathbf{c})|\leq L\exp(-K),\quad|G_{h}(x;\mathbf{c})-G_{h_{j}}(x;\mathbf{c})|\leq L\exp(-K),$ (3.10)
where $L$ is a Lipschitz constant depending on $R$. Thus, if
$\min(|F_{h_{j}}(x;\mathbf{c})|,|G_{h_{j}}(x;\mathbf{c})|)\geq 2L\exp(-K),$
then $\min(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|)\geq L\exp(-K)$. This allows us to transfer the Cartan estimate from $h_{j}$ to all $h$ in its neighborhood. Specifically, for each $h_{j}$, we have from Step (2) that
$\text{mes}\left(\mathcal{B}^{(iii)}(h_{j},K)\cap B_{R}\right)=0\quad\text{for }K\geq K_{0}(R,h_{j}).$ (3.11)
By the covering, the union over $h$ can be controlled by the sum over $j$:
$\text{mes}\left(\bigcup_{h\in\mathbb{T}^{d}}\mathcal{B}^{(iii)}(h,K)\cap B_{R}\right)\leq\sum_{j=1}^{N(K)}\text{mes}\left(\mathcal{B}^{(iii)}(h_{j},K)\cap B_{R}\right).$ (3.12)
From (3.6) and the choice of $K_{0}$, for $K\geq\max_{j}K_{0}(R,h_{j})$, each term in the sum is zero. For $K<\max_{j}K_{0}(R,h_{j})$, we use a coarse bound: since $\mathcal{B}^{(iii)}(h_{j},K)\subset B_{R}$ and $\text{mes}(B_{R})<\infty$, we have
$\text{mes}\left(\bigcup_{h\in\mathbb{T}^{d}}\mathcal{B}^{(iii)}(h,K)\cap B_{R}\right)\leq N(K)\cdot\text{mes}(B_{R})\leq M\exp(Kd)\cdot\exp(-C^{\prime}K)$ (3.13)
for some $M,C^{\prime}>0$, where $C^{\prime}$ is chosen such that $\exp(-C^{\prime}K)$ dominates the measure of the bad set for each $h_{j}$. Since
$\operatorname{mes}(\mathcal{B}^{(iii)}(h_{j},K)\cap B_{R})\leq\exp(-C^{\prime}K)$
for some $C^{\prime}>d$ (from the Cartan estimate), the series
$\sum_{K\geq K_{0}}\operatorname{mes}(\mathcal{B}^{(iii)}(h_{j},K)\cap B_{R})$
converges. By the Borel-Cantelli lemma, the set
$\bigcup_{K\geq K_{0}}\bigcup_{h\in\mathbb{T}^{d}}\mathcal{B}^{(iii)}(h,K)\cap B_{R}$
has measure zero. Since $\mathbb{R}^{N}=\bigcup_{R\in\mathbb{N}}B_{R}$ and measure zero is preserved under countable unions, the full bad set $\mathcal{B}^{(iii)}$ has measure zero.
We now turn to the genericity of the gradient Cartan estimate (condition (iv)). Let $\mathcal{B}^{(iv)}$ be the set of all coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which the gradient Cartan estimate fails, i.e., there exists some $\eta\in\mathbb{R}$, some unit vector $h_{0}\in\mathbb{R}^{d}$, and some $K\geq K_{0}$ such that:
$\text{mes}\left\{x\in\mathbb{T}^{d}:\min\left(|V(x;\mathbf{c})-\eta|,|\langle\nabla V(x;\mathbf{c}),h_{0}\rangle|\right)<\exp(-K)\right\}>\exp(-K^{\mathfrak{c}_{1}}).$
More formally, we can express this as:
$\mathcal{B}^{(iv)}=\bigcup_{h_{0}\in S^{d-1}}\bigcup_{\eta\in\mathbb{R}}\bigcup_{K\geq K_{0}}\mathcal{B}^{(iv)}(\eta,h_{0},K)$
where $S^{d-1}$ is the unit sphere in $\mathbb{R}^{d}$, and for fixed $\eta$, $h_{0}$, and $K$, the set $\mathcal{B}^{(iv)}(\eta,h_{0},K)$ is defined as:
$\mathcal{B}^{(iv)}(\eta,h_{0},K)$ (3.14)
$=\left\{\mathbf{c}\in\mathbb{R}^{N}:\text{mes}\left\{x:\min\left(|V(x;\mathbf{c})-\eta|,|\langle\nabla V(x;\mathbf{c}),h_{0}\rangle|\right)<\exp(-K)\right\}>\exp(-K^{\mathfrak{c}_{1}})\right\}.$ (3.15)
The argument for showing that $\mathcal{B}^{(iv)}$ has measure zero parallels that for $\mathcal{B}^{(iii)}$ but requires careful treatment of the additional parameters $\eta$ and $h_{0}$.
For the analysis, it is convenient to define the relevant analytic functions:
$F_{\eta}(x;\mathbf{c})=V(x;\mathbf{c})-\eta,\quad G_{h_{0}}(x;\mathbf{c})=\langle\nabla V(x;\mathbf{c}),h_{0}\rangle.$ (3.16)
In terms of these functions, the gradient Cartan estimate requires that for any $K\gg 1$,
$\text{mes}\left\{x\in\mathbb{T}^{d}:\min\left(|F_{\eta}(x;\mathbf{c})|,|G_{h_{0}}(x;\mathbf{c})|\right)<\exp(-K)\right\}\leq\exp(-K^{\mathfrak{c}_{1}}).$ (3.17)
The corresponding bad set for fixed $\eta$, $h_{0}$, and $K$ is:
$\mathcal{B}^{(iv)}(\eta,h_{0},K)$ (3.18)
$=\left\{\mathbf{c}\in\mathbb{R}^{N}:\text{mes}\left\{x:\min\left(|F_{\eta}(x;\mathbf{c})|,|G_{h_{0}}(x;\mathbf{c})|\right)<\exp(-K)\right\}>\exp(-K^{\mathfrak{c}_{1}})\right\}.$ (3.19)
The overall bad set is
$\mathcal{B}^{(iv)}=\bigcup_{h_{0}\in S^{d-1}}\bigcup_{\eta\in\mathbb{R}}\bigcup_{K\geq K_{0}}\mathcal{B}^{(iv)}(\eta,h_{0},K),$
where $S^{d-1}$ is the unit sphere.
Since $\sum_{K\geq K_{0}}\exp(-(C^{\prime}-d)K)<\infty$ when $C^{\prime}>d$, by the Borel–Cantelli lemma, the union of the bad sets has measure zero.
The proof proceeds using the same three-step structure as for $\mathcal{B}^{(iii)}$:
(4) Uniform Estimate for Bounded Parameters: For a fixed radius $R>0$ and parameters $\mathbf{c}$ in the ball $B_{R}=\{\mathbf{c}\in\mathbb{R}^{N}:\|\mathbf{c}\|\leq R\}$, the functions $F_{\eta}$ and $G_{h_{0}}$ are analytic in $x$ and linear in $\mathbf{c}$. Applying the Cartan-type estimate yields constants $\tilde{C}_{R},\tilde{\delta}_{R}>0$ such that for $\mathbf{c}\in B_{R}$ (excluding a set where $F_{\eta}$ and $G_{h_{0}}$ are identically zero, which has measure zero),
$\text{mes}\left\{x:\min\left(|F_{\eta}(x;\mathbf{c})|,|G_{h_{0}}(x;\mathbf{c})|\right)<\epsilon\right\}\leq\tilde{C}_{R}\epsilon^{\tilde{\delta}_{R}}.$ (3.20)
(5) Measure Zero for Bounded Parameters: For fixed $\eta$, $h_{0}$, and $R$, we choose $K_{0}(R,\eta,h_{0})$ large enough so that for $K\geq K_{0}$,
$\tilde{C}_{R}\exp(-\tilde{\delta}_{R}K)\leq\exp(-K^{\mathfrak{c}_{1}}).$
Then, for such $K$,
$\mathcal{B}^{(iv)}(\eta,h_{0},K)\cap B_{R}=\emptyset.$
For $K<K_{0}$, the bad set is contained in the null set where the functions are identically zero. Thus, $\bigcup_{K\geq K_{0}}\mathcal{B}^{(iv)}(\eta,h_{0},K)\cap B_{R}$ has measure zero.
(6) Union over $\eta$, $h_{0}$, $K$, and Unbounded Parameters: This is the most critical step, addressing the union over the non-compact set $\mathbb{R}$ for $\eta$ and the compact set $S^{d-1}$ for $h_{0}$.
  • Union over $\eta$: Although $\eta\in\mathbb{R}$ is non-compact, the function $V(x;\mathbf{c})$ is bounded for $\mathbf{c}\in B_{R}$. Let $I_{R}$ be a compact interval containing the range of $V$ for all $\mathbf{c}\in B_{R}$, e.g.,
    $I_{R}=[-\max_{\mathbf{c}\in B_{R}}\|V(\cdot;\mathbf{c})\|_{\infty},\max_{\textbf{c}\in B_{R}}\|V(\cdot;\textbf{ c})\|_{\infty}].$
    If $\eta\notin I_{R}$, then $|V(x;\mathbf{c})-\eta|$ is bounded away from zero uniformly, and the Cartan estimate holds trivially. Therefore, it suffices to consider $\eta\in I_{R}$, a compact set. We can cover $I_{R}$ with $O(\exp(K))$ intervals of length $\exp(-K)$.
  • Union over $h_{0}$: The unit sphere $S^{d-1}$ is compact. We cover it with a grid of $O(\exp(K(d-1)))$ points $h_{0,j}$ with mesh size $\exp(-K)$.
  • Using the uniform continuity of the functions with respect to $\eta$ and $h_{0}$, the estimate for a fixed $(\eta_{j},h_{0,k})$ can be extended to a neighborhood of size $\exp(-K)$ around them, with a possible adjustment of constants.
Following the same covering argument and application of the Borel-Cantelli lemma as in Step (3) of the original proof, and then taking the union over $R\in\mathbb{N}$, we conclude that the full set $\mathcal{B}^{(iv)}$ has Lebesgue measure zero.
Therefore, the set of coefficients for which any Cartan estimate fails has Lebesgue measure zero. ∎
Proof of Theorem 2.2.
The theorem asserts that the set of coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which $V(\cdot;\mathbf{c})\in\mathfrak{G}$ has full Lebesgue measure. Recall that $\mathfrak{G}$ is defined by four conditions (Definition 2.1). Let $\mathcal{B}=\mathcal{B}^{(i)}\cup\mathcal{B}^{(ii)}\cup\mathcal{B}^{(iii)}\cup\mathcal{B}^{(iv)}$ be the set of parameters where at least one condition fails.
By Lemma 3.1, the sets $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$ have measure zero. By Lemma 3.2, the sets $\mathcal{B}^{(iii)}$ and $\mathcal{B}^{(iv)}$ also have measure zero. Since a finite union of measure zero sets has measure zero, $\mathcal{B}$ has measure zero. Consequently, the set of good parameters $\mathcal{G}=\mathbb{R}^{N}\setminus\mathcal{B}$ has full Lebesgue measure. This completes the proof of the genericity conjecture by Goldstein et al. [11]. ∎
@@
The proof of Theorem 2.2 actually establishes a stronger result: the set of good parameters $\mathcal{G}$ is not only of full measure but also a residual set in $\mathbb{R}^{N}$. This follows because the exceptional sets $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$, as shown in Lemma 3.1, are contained in smooth submanifolds of positive codimension, which are closed and have empty interior (hence nowhere dense). The exceptional sets $\mathcal{B}^{(iii)}$ and $\mathcal{B}^{(iv)}$ from Lemma 3.2 can also be shown to be of first category. Thus, $\mathcal{G}$ is the complement of a finite union of first category sets, making it a residual set. This confirms that the potentials in class $\mathfrak{G}$ are generic in both the measure-theoretic and topological senses.
@@ Acknowledgments
This work was supported by NSFC (No. 11571327, 11971059)

### v2

@@ 3.2 Genericity of Cartan Estimates
We first treat condition (iii). For each $R>0$, let $\mathcal{Z}_{R}$ be the null set obtained from Proposition 3.5 (which depends on $R$ because the constants $C_{R},\gamma_{R}$ in Proposition 3.2 depend on $R$). Actually Proposition 3.5 was proved for a fixed $R$, so we denote by $\mathcal{Z}_{R}$ the null set such that for every $\mathbf{c}\notin\mathcal{Z}_{R}$ and every $h\in\mathbb{T}^{d}$, condition (iii) holds. Set $\mathcal{Z}^{(iii)}=\bigcup_{R\in\mathbb{N}}\mathcal{Z}_{R}$; then $\mathcal{Z}^{(iii)}$ is a countable union of null sets, hence null. For any $\mathbf{c}\notin\mathcal{Z}^{(iii)}$, there exists $R$ with $\|\mathbf{c}\|\leq R$, so $\mathbf{c}\notin\mathcal{Z}_{R}$ and condition (iii) holds for all $h$.
Similarly, for condition (iv) we obtain a null set $\mathcal{Z}^{(iv)}$ (by taking the union over $R$ of the null sets from the gradient version of Proposition 3.5). Finally, let $\mathcal{Z}=\mathcal{Z}^{(iii)}\cup\mathcal{Z}^{(iv)}$. Then $\mathcal{Z}$ is null, and for every $\mathbf{c}\notin\mathcal{Z}$ both Cartan estimates (iii) and (iv) are satisfied. Therefore the set of coefficients for which the estimates fail is contained in $\mathcal{Z}$ and has Lebesgue measure zero. ∎
Proof of Theorem $2.$
Recall that $\mathfrak{G}$ is defined by four conditions (Definition 2.1). Let $\mathcal{B}=\mathcal{B}^{(i)}\cup\mathcal{B}^{(ii)}\cup\mathcal{B}^{(iii)}\cup\mathcal{B}^{(iv)}$ be the set of parameters where at least one condition fails. By Lemma 3.1, $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$ have measure zero. By Lemma 3.6, $\mathcal{B}^{(iii)}$ and $\mathcal{B}^{(iv)}$ also have measure zero. A finite union of measure-zero sets has measure zero, so $\mathcal{B}$ is null. Hence the set of good parameters $\mathcal{G}=\mathbb{R}^{N}\setminus\mathcal{B}$ has full Lebesgue measure. This proves the genericity conjecture of Goldstein, Schlag, and Voda [13]. ∎
@@
The proof of Theorem 2.2 actually establishes a stronger result: the set of good parameters $\mathcal{G}$ is not only of full measure but also a residual set in $\mathbb{R}^{N}$. This follows because the exceptional sets $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$ are contained in smooth submanifolds of positive codimension, which are closed and have empty interior (hence nowhere dense). The exceptional sets $\mathcal{B}^{(iii)}$ and $\mathcal{B}^{(iv)}$ can also be shown to be of first category. Thus $\mathcal{G}$ is the complement of a finite union of first category sets, making it residual. This confirms that the potentials in class $\mathfrak{G}$ are generic in both the measure-theoretic and topological senses.
@@ Acknowledgments
This work was supported by NSFC (No. 11571327, 11971059).